BUSI 320 Comprehensive Problem 2 Spring 2019

You have been asked to assess the expected financial impact of each of the following proposals to improve the profitability of credit sales made by your company. Each proposal is independent of the other. Answer all questions. Showing your work may earn you partial credit.

Proposal #1 would extend trade credit to some customers that previously have been denied credit because they were considered poor risks.   Sales are projected to increase by $120,000 per year if credit is extended to these new customers. Of the new accounts receivable generated, 6% are projected to be uncollectible. Additional collection costs are projected to be 5% of incremental sales, and production and selling costs are projected to be 80% of sales. Your firm expects to pay a total of 30% of its income after expenses in taxes.

  1. Compute the incremental income after taxes that would result from these projections:
  1. Compute the incremental Return on Sales if these new credit customers are accepted:

If the receivable turnover ratio is expected to be 4 to 1 and no other asset buildup is needed to serve the new customers…

  1. Compute the additional investment in Accounts Receivable
  2. Compute the incremental Return on New Investment
  1. If your company requires a 20% Rate of Return on Investment for all proposals, do the numbers suggest that trade credit should be extended to these new customers? Explain.

Proposal #2 would establish local collection centers throughout the region to decrease the time it takes to convert credit payments that are mailed in by check to cash. It is estimated that establishing these collection centers would reduce the average collection time by 2 days.

  1. If the company currently averages $25,000 in collections per day, how many dollars will this suggested cash management system frees up?

  1. If all freed up dollars would be used to pay down debt that has an interest rate of 8%, how much money could be saved each year in interest expense?
  1. Do the numbers suggest that this new system should be implemented if its total annual cost is $7200? Explain.
Answer #1

Proposal 1

Part (a)

Compute the incremental income after taxes that would result from these projections:

Please see the table below. The column "Linkage" explains how each row has been calculated. The last line in the table is the answer.

Incremental P&L Linkage $
Incremental Sales S           120,000
[-] Uncollectibles 6% x S               7,200
[-] Additional collection costs 5% of S               6,000
[-] production & selling costs 80% of S             96,000
Incremental profit before taxes             10,800
[-] Taxes 30% of Incremental profit               3,240
Incremental income after taxes               7,560

Part (b)

Compute the incremental Return on Sales if these new credit customers are accepted:

Incremental Return on Sales = Incremental net income / Incremental sales = 7,560 / 120,000 = 6.30%

Part (c)

Compute the additional investment in Accounts Receivable

Receivable turnover = Incremental Sales/ Incremental receivables = 4

Hence, incremental receivables = incremental sales / 4 = 120,000 / 4 = 30,000

Part (d) Compute the incremental Return on New Investment

Incremental return on new investment = incremental net income / incremental investment in receivables = 7,560 / 30,000 = 25.20%

Part (e) If your company requires a 20% Rate of Return on Investment for all proposals, do the numbers suggest that trade credit should be extended to these new customers? Explain.

Since, incremental return on trade credit as calculated above = 25.20% > 20% = company's required return on investment , the numbers do suggest that trade credit should be extended to these new customers.

Proposal 2

If the company currently averages $25,000 in collections per day, how many dollars will this suggested cash management system frees up?

Average collection per day = $ 25,000

Reduction in collection period = 2 days

So, the amount freed up by this cash management system =  Average collection per day x Reduction in collection period = = $ 25,000 x 2 days = $ 50,000

If all freed up dollars would be used to pay down debt that has an interest rate of 8%, how much money could be saved each year in interest expense?

Money that could be saved = interest on $ 50,000 @ 8% = 50,000 x 8% = $ 4,000

Do the numbers suggest that this new system should be implemented if its total annual cost is $7200? Explain.

Since the annual benefit of the new system = $ 4,000 < $ 7,200 = annual cost of implementation, the numbers don't suggest implementation of this system.